OpenStudy (anonymous):

Are 24, 45, and 51 the side measurements of a special right triangle? Yes or No, Explain your answer and show work.

OpenStudy (anonymous):

@Opcode

OpenStudy (anonymous):

24^2 =576, 45^2=2025, 51^2 =2601 and 2601=2025+576. so according to Pythagorean theorem, those form a right triangle.

OpenStudy (anonymous):